OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (lastdaywork):

Simply use the fact that - 3x + 2 = 14

OpenStudy (anonymous):

is that -3x or....

OpenStudy (lastdaywork):

3x + 2 = 14 :)

OpenStudy (anonymous):

is that it

OpenStudy (lastdaywork):

Yep! Were you looking for something complex ?? :P

OpenStudy (anonymous):

YEA (0.0)

OpenStudy (lastdaywork):

In that case, here we go - Let there be a function g defined as \[g(y) = f^{-1}(y)\] Then \[g(f(x)) = f^{-1}(f(x))=x\] We were given that - f(x) = 3x + 2 implies \[x=\frac{ f(x) - 2 }{ 3 }\] \[g(f(x)) = \frac{ f(x) - 2 }{ 3 }\] \[g(y)=\frac{ y-2 }{ 3 }\] The question said that f(x) = 14 implies x = g(14) Now can you find g(14) ?

OpenStudy (lastdaywork):

Note that if f is not bijective; then g will not be a function in the first place. Or in other words, we may not get a unique value of x satisfying f(x) = 14

OpenStudy (anonymous):

so this inst the rite choice Given that f(x) = 14, we substitute 14 in to the equation in place of f(x) 14 = 3x + 2 Subtract two from both sides of the equation 12 = 3x Divide both sides of the equation by three. 4 = x Answer: 4

OpenStudy (lastdaywork):

"so this inst the rite choice" IDK what does that mean.. Yea, you can (& should) simply solve 14 = 3x + 2 I only tried to justify this ^^ step using the concept of inverse of functions. ;)

OpenStudy (anonymous):

srry (isn't) and thnx

OpenStudy (lastdaywork):

You're welcome :)